Mathematics 243 – Algebraic Structures Definitions Quiz 2 – September 14, 2012

Your Name:

A) (5) Let f : A → B be a mapping. What does it mean to say that f is one-to-one? (Any correct and precisely stated form is OK; only one form is necessary. Don’t write too much!)

B) (5) Let ∗ be the binary operation on Z defined by x ∗ y = 4x − y. Is ∗ associative? Why or why not?